# Non-thermal production of Higgsino dark matter by late-decaying scalar fields

### v1

@@ Abstract
We study the non-thermal production of the Higgsino dark matter (DM). Assuming that the lightest neutral Higgsino is the lightest supersymmetric particle (LSP) in minimal supersymmetric standard model, we calculate the relic abundance of the Higgsino LSP produced by the decay of late-decaying scalar field. In the calculation of the relic abundance, we have properly included the effects of coannihilation as well as the non-perturbative effect (known as the Sommerfeld effect). Contrary to the case of the thermal-relic scenario, in which the observed DM abundance is realized with the Higgsino mass of $\sim 1.2\ {\rm TeV}$, Higgsino DM is possible with lighter Higgsino mass as the reheating temperature becomes lower than the Higgsino mass. The reheating temperature relevant for realizing the correct DM density is presented as a funciton of the Higgsino mass.
@@ I Introduction
In this paper, we study the non-thermal DM production scenario. In order to precisely calculate the relic abundance of the Higgsino DM in the non-thermal production scenario, we should take into account several effects which may complicate the calculation of the relic abundance. First, there exist three mass eigenstates, i.e., the lightest neutral Higgsino (i.e., the DM candidate) $\chi^{0}_{1}$, second-lightest neutral Higgsino $\chi^{0}_{2}$, and the chargid Higgsino $\chi^{\pm}_{1}$, thus the coannihilations among these significantly affect the relic abundance. In addition, because Higgsinos have electroweak quantum numbers, the Sommerfeld effect may enhance the annihilation cross sections of Higgsinos Hisano et al. (2004, 2005, 2007). The mass splittings among the mass eigenstates, which are model dependent, may also affect the relic abundance. In this paper, we carefully take into account these effects and calculate the relic density of the lightest neutral Higgsino. Although the non-thermal production of the Higgsino DM was considered before Han (2019), these effects (particularly, the Sommerfeld effect) have not been fully considered. Thus, in this paper, we carefully taken into account these issues and examine the production process of the Higgsino DM from the decay of long-lived scalar field.
@@ II Higgsino DM
$\frac{\partial m_{\text{LSP}}}{\partial v}=\frac{\Delta m_{0}}{v}\quantity(1+\sigma\sin 2\beta),$ (15)
@@
Because the mass splittings among Higgsinos are model dependent, as explained so far, we consider several cases in our numerical calculation. Given the experimental constraints discussed above, the mass splittings we adopt are summarised in Table 1.
@@ III Non-thermal production of DM
where $\kappa$ is a free parameter and $M_{\rm Pl}\simeq 2.4\times 10^{18}\ {\rm GeV}$ is the reduced Planck scale.
@@
where $s_{\rm rad}$ is the entropy density of radiation, $\langle E_{\rm DM}\rangle$ (which is taken to be $\sqrt{m_{\rm DM}^{2}+T^{2}}$ in our numerical calculation) is the thermally averaged energy of DM particle, and
@@
where $s(T)$ is the entropy density of the universe as a function of the cosmic temperature $T$, and $T_{\rm now}\simeq 2.7\,{\rm K}$ is the present cosmic temperature.
Although the above argument provides an order-of-magnitude estimation of the DM relic density, more careful analysis is needed for a precise prediction of the DM abundance. For this purpose, in the following, we numerically solve the Boltzmann equations numerically and calculate the relic abundance of the Higgsino DM for various choices of model parameters.
@@ IV.1 Sommerfeld effects
Sommerfeld effect is a non-perturbative quantum effect caused by exchange of light mediators. Particularly, for the study of the DM relic abundance, the Sommerfeld effect may significantly enhance or suppresse the annihilation cross section of DM pairs so it may play important role Hisano et al. (2004, 2005, 2007). For the case of DM candidates having electroweak quantum numbers, like the wino or Higgsino in the supersymmetric model or $SU(2)_{L}$ quintuplet, the Sommerfeld effect is known to sigificantly alter the thermal relic abundance. Because the Higgsino multiplet contains several mass eigenstates which have relatively small mass splittings, as explained in section II, we should take into account the coannihilations of neutralinos $\chi_{1}^{0}$, $\chi_{2}^{0}$ and charginos $\chi_{1}^{\pm}$.
In order to study the Sommerfeld effect on the annihilation rates, we derive the effective actions for the non-relativistic two-body states built from pairs of $\chi_{1}^{0}$, $\chi_{2}^{0}$ and $\chi_{1}^{\pm}$. Following the procedure in Ref. Hisano et al. (2005), we firstly integrate out the gauge bosons in the MSSM which give the interaction action between currents of Higgsinos. Then, the high momentum modes are integrated out, inducing the absorptive parts in the non-relativestic effective action. The final effective action has the form of
@@
where $\Phi$ are vectors of composite field, the argument $x$ is the center-of-mass coordinate of the two-body states and $\vec{r}$ is the relative coordinate vector. The components of $\Phi$ are built from two-body states with the same spin $S$ and charge $Q$. The potential $\mathcal{H}(r)$ describes the Coulomb and Yukawa forces given by the exchange of gauge bosons. The absorptive part $\Gamma$ corresponds to short-distance annihilations of two-body states, can be obtained by calculating diagrams of transitions between two-body states and using the optical theorem. We calculate $\Gamma$ with taking into account all the relevant box, triangle, and self-energy diagrams. The formulas for $\mathcal{H}(r)$ and $\Gamma$ are shown in the appendix.
@@
$\sigma_{i\rightarrow\rm{light}}v_{\rm{rel}}$ $=\sum_{S}c_{i}\times s_{i}\times\Psi_{ij}^{\infty}(\Gamma_{Q}^{S})_{jk}\Psi^{\infty*}_{ki}$ (45)
@@ IV.1.1 Sommerfeld enhancement
In order to show the typical behavior of the Sommerfeld factors, in Fig. 1, we plot the Sommerfeld factors for $Q=0$ states $\chi_{1}^{0}\chi_{1}^{0}$, $\chi_{2}^{0}\chi_{2}^{0}$ and $\chi^{+}\chi^{-}$ through the ${}^{1}S_{0}$ partial-wave annihilations as functions of $v_{*}$. Here, we take $m_{\chi_{0}^{1}}=1.1\ {\rm TeV}$ and adopt the mass splittings of Model A given in Table 1 (i.e., $\Delta m_{\chi^{\pm}}=0.35\ {\rm GeV}$ and $\Delta m_{\chi^{0}}=0.1\ {\rm GeV}$), for which the channel $\chi_{2}^{0}\chi_{2}^{0}$ opens at $v_{*}\simeq 0.02697$, and the threshold for appearance of $\chi^{+}\chi^{-}$ in the asymptotic state is $v_{*}\simeq 0.0505$. At small velocities, the Sommerfeld factor approaches to $\sim 1.715$ and velocity-independent, which is a typical behavior given by Yukawa potential. On the other hand, at large velocity, the perturbative picture becomes relevant and the Sommerfeld factor approaches to $1$. The behavior of the Sommerfeld factor near the threshold is rather complicated. Just above the threshold of $\chi^{+}\chi^{-}$, the on-shell charged Higgsinos state can be produced and the enhancement factor of $\chi^{+}\chi^{-}$ diverges as the inverse of relative velocity between them $1/v^{\chi^{\pm}}_{\rm{rel}}$. Such divergence is expected because at large distance, the Coulomb potential dominates the interactions. In this regime, an analytic estimate to the Sommerfeld factor of $\chi^{+}\chi^{-}$ is given by Landau and Lifshits (1991)
@@
Figure 1: The Sommerfeld enhancement factors for the ${}^{1}S_{0}$ partial-wave annihilations of $\chi_{1}^{0}\chi_{1}^{0}$, $\chi_{2}^{0}\chi_{2}^{0}$ and $\chi^{+}\chi^{-}$ channels. $v_{*}$ is the relative velocity between $\chi_{1}^{0}$s in the lightest two-particle state $\chi_{1}^{0}\chi_{1}^{0}$. For other heavier channels, the relative velocity $v_{\rm{rel}}$ is related to $v_{*}$ by the mass splittings.
@@ IV.1.2 Thermally averaged cross sections
$g_{\rm{eff}}=2+2(1+\Delta m_{\chi^{0}}/m_{\chi_{1}^{0}})^{3/2}e^{-x\Delta m_{\chi^{0}}/m_{\chi_{1}^{0}}}+4(1+\Delta m_{\chi^{\pm}}/m_{\chi_{1}^{0}})^{3/2}e^{-x\Delta m_{\chi^{\pm}}/m_{\chi_{1}^{0}}}.$ (50)
@@
In Fig. 3, we plot $\langle\sigma v_{\rm{rel}}\rangle$ as a function of $x$ with and without including the Sommerfeld effect. For the chosen values of mass splittings, the decoupling of heavier neutralino and charginos occurs at around $x\sim 10^{4}$. After the decoupling, the enhancement of $\langle\sigma v_{\rm{rel}}\rangle$ becomes constant, due to the velocity-independent Sommerfeld factor of $\chi_{1}^{0}\chi_{1}^{0}$ at very low velocities as shown in Fig. 1.
Figure 3: The perturbative and Sommerfeld enhanced thermally averaged cross section $\langle\sigma v_{\rm{rel}}\rangle$ versus $x=m_{\chi_{1}^{0}}/T$.
@@ IV.2 Relic abundances
$\frac{dT}{dA}$ $=\left(1+\frac{T}{3h_{*}}\frac{dh_{*}}{dT}\right)^{-1}\biggl{[}-\frac{T}{A}\frac{15T_{R}^{6}}{2\pi^{2}\sqrt{3}M_{\rm{Pl}}HT^{3}h_{*}A^{11/2}}\biggl{(}c_{\rho}^{1/2}A^{3/2}(1-B)\Phi$
@@ VI Conclusions and discussion
We have paid particular attention to the Higgsino LSP scenario, which is one of the popular and well-motivated scenario in the MSSM. We have performed a detailed analysis of the non-thermal production processes of Higgsino DM due to the decay of late-decaying scalar field. We have assumed that superparticles other than Higgsinos are so heavy that they do not affect the relic abundance of the Higgsino DM via coannihilation. We have numerically solved the Boltzmann equation giverning the number density of Higgsino LSP, taking into accout all the relevant annihilating channels among Higgsinos. Because Higgsinos have gauge quantum numbers and couple to electroweak gauge bosons, a non-perturbative effect, so called the Sommerfeld effect, may be sigificant; in our calculation, the Sommerfeld effect has been properly taken into account in the calculation of the annihilation cross sections of Higgsinos.
@@ A.0.2 $Q=0,S=1$
$\mathcal{H}_{Q=0}^{S=1}=\Phi_{0}^{V\dagger}\begin{pmatrix}2\Delta m_{\chi^{0}}-\Delta m_{\chi^{\pm}}-\frac{A^{2}\alpha_{2}e^{-m_{Z}r}}{4c_{w}^{2}r}-\frac{\alpha}{r}&-i\frac{\alpha_{2}e^{-m_{W}r}}{4r}\\ i\frac{\alpha_{2}e^{-m_{W}r}}{4r}&-\frac{\alpha_{2}e^{-m_{Z}r}}{4c_{w}^{2}r}\end{pmatrix}\Phi_{0}^{V}.$ (63)

### v3

@@ Abstract
We study the non-thermal production of the Higgsino dark matter (DM). Assuming that the lightest neutral Higgsino is the lightest supersymmetric particle (LSP) in the minimal supersymmetric standard model, we calculate the relic abundance of the Higgsino LSP produced by the decay of late-decaying scalar field. In the calculation of the relic abundance, we have properly included the effects of coannihilation as well as the non-perturbative effect (known as the Sommerfeld effect). Contrary to the case of the thermal-relic scenario, in which the observed DM abundance is realized with the Higgsino mass of $\sim 1.2\ {\rm TeV}$, Higgsino DM is possible with a lighter Higgsino mass as the reheating temperature becomes lower than the Higgsino mass. The reheating temperature relevant for realizing the correct DM density is presented as a function of the Higgsino mass.
@@ I Introduction
In this paper, we study the non-thermal DM production scenario. In order to precisely calculate the relic abundance of the Higgsino DM in the non-thermal production scenario, we should take into account several effects which may complicate the calculation of the relic abundance. First, there exist three mass eigenstates, i.e., the lightest neutral Higgsino (i.e., the DM candidate) $\chi^{0}_{1}$, second-lightest neutral Higgsino $\chi^{0}_{2}$, and the charged Higgsino $\chi^{\pm}_{1}$, thus the coannihilations among these significantly affect the relic abundance. In addition, because Higgsinos have electroweak quantum numbers, the Sommerfeld effect may enhance the annihilation cross sections of Higgsinos Hisano et al. (2004, 2005, 2007). The mass splittings among the mass eigenstates, which are model dependent, may also affect the relic abundance. In this paper, we carefully take into account these effects and calculate the relic density of the lightest neutral Higgsino. Although the non-thermal production of the Higgsino DM was considered before Han (2019), these effects (particularly, the Sommerfeld effect) have not been fully considered. Thus, in this paper, we carefully taken into account these issues and examine the production process of the Higgsino DM from the decay of long-lived scalar field.
@@ II Higgsino DM
$\frac{\partial m_{\text{LSP}}}{\partial v}=\frac{\Delta m_{\chi^{0}}}{v}\quantity(1+\sigma\sin 2\beta),$ (15)
@@
Because the mass splittings among Higgsinos are model dependent, as explained so far, we consider several cases in our numerical calculation. Given the experimental constraints discussed above, the mass splittings we adopt are summarized in Table 1.
@@ III Non-thermal production of DM
where $\kappa$ is a free parameter and $M_{\rm Pl}\simeq 2.4\times 10^{18}\ {\rm GeV}$ is the reduced Planck scale. For the moduli field in string theory, $\kappa\sim O(1)$ is expected. For other long-lived fields, such as the saxion, the cut-off scale giving rise to the dimension-5 operator given in Eq. (22) may be lower than the reduced Planck scale, which corresponds to $\kappa\gg O(1)$. Our following analysis is applicable to both cases with long-lived scalar fields which once dominate the universe and we do not exclude the latter possibility.
@@
where $s_{\rm rad}$ is the entropy density of radiation, $\langle E_{\rm DM}\rangle$ (which is taken to be $m_{\rm DM}+\frac{3}{2}T$ in our numerical calculation) is the thermally averaged energy of DM particle, and
@@
where $s(T)$ is the entropy density of the universe as a function of the cosmic temperature $T$, and $T_{\rm now}\simeq 2.7\,{\rm K}$ is the present cosmic temperature. Using the relation between the reheating temperature $T_{\rm R}$ and the decay rate, we can see that, when $B_{\rm DM}\gtrsim B_{*}$, the annihilation of the DM becomes so efficient at the time of the reheating that the first term in the right-hand side of Eq. (38) dominates over the second term; using the relation between the reheating temperature $T_{\rm R}$ and the decay rate given in Eq. (27), $B_{*}$ is estimated as
$B_{*}\sim\frac{m_{\phi}}{M_{\rm Pl}T_{\rm R}^{2}\langle\sigma v\rangle}\sim 10^{-5}\times\left(\frac{m_{\phi}}{100\ {\rm TeV}}\right)\left(\frac{m_{\rm DM}}{1\ {\rm TeV}}\right)^{2}\left(\frac{T_{\rm R}}{1\ {\rm GeV}}\right)^{-2},$ (39)
where, in the second equality, we used the estimation of the Higgsino annihilation cross section.
Although the above argument provides an order-of-magnitude estimation of the DM relic density, more careful analysis is needed for a precise prediction of the DM abundance. For this purpose, in the following, we solve the Boltzmann equations numerically and calculate the relic abundance of the Higgsino DM for various choices of model parameters.
@@ IV.1 Sommerfeld effects
Sommerfeld effect is a non-perturbative quantum effect caused by exchange of light mediators. Particularly, for the study of the DM relic abundance, the Sommerfeld effect may significantly enhance or suppress the annihilation cross section of DM pairs so it may play an important role Hisano et al. (2004, 2005, 2007). For the case of DM candidates having electroweak quantum numbers, like the wino or Higgsino in the supersymmetric model or $SU(2)_{L}$ quintuplet, the Sommerfeld effect is known to significantly alter the thermal relic abundance. Because the Higgsino multiplet contains several mass eigenstates which have relatively small mass splittings, as explained in section II, we should take into account the coannihilations of neutralinos $\chi_{1}^{0}$, $\chi_{2}^{0}$ and charginos $\chi_{1}^{\pm}$.
In order to study the Sommerfeld effect on the annihilation rates, we derive the effective actions for the non-relativistic two-body states built from pairs of $\chi_{1}^{0}$, $\chi_{2}^{0}$ and $\chi_{1}^{\pm}$. Following the procedure in Ref. Hisano et al. (2005), we firstly integrate out the gauge bosons in the MSSM which give the interaction action between currents of Higgsinos. Then, the high momentum modes are integrated out, inducing the absorptive parts in the non-relativistic effective action. The final effective action has the form of
@@
where $\Phi$ are vectors of composite field, the argument $x$ is the center-of-mass coordinate of the two-body states and $\vec{r}$ is the relative coordinate vector. The components of $\Phi$ are built from two-body states with the same spin $S$ and charge $Q$. The potential $\mathcal{H}(r)$ describes the Coulomb and Yukawa forces given by the exchange of gauge bosons. The absorptive part $\Gamma$ corresponds to short-distance annihilation of two-body states, can be obtained by calculating diagrams of transitions between two-body states and using the optical theorem. We calculate $\Gamma$ with taking into account all the relevant box, triangle, and self-energy diagrams. The formulas for $\mathcal{H}(r)$ and $\Gamma$ are shown in the appendix.
@@
$\sigma_{i\rightarrow\rm{light}}v_{\rm{rel}}$ $=\sum_{S}c_{i}\times s_{i}\times\Psi_{ij}^{\infty}(\Gamma_{Q}^{S})_{jk}\Psi^{\infty*}_{ik}$ (46)
@@ IV.1.1 Sommerfeld enhancement
In order to show the typical behavior of the Sommerfeld factors, in Fig. 1, we plot the Sommerfeld factors for $Q=0$ states $\chi_{1}^{0}\chi_{1}^{0}$, $\chi_{2}^{0}\chi_{2}^{0}$ and $\chi^{+}\chi^{-}$ through the ${}^{1}S_{0}$ partial-wave annihilation as functions of $v_{*}$. Here, we take $m_{\chi_{0}^{1}}=1.1\ {\rm TeV}$ and adopt the mass splittings of Model A given in Table 1 (i.e., $\Delta m_{\chi^{\pm}}=0.35\ {\rm GeV}$ and $\Delta m_{\chi^{0}}=0.1\ {\rm GeV}$), for which the channel $\chi_{2}^{0}\chi_{2}^{0}$ opens at $v_{*}\simeq 0.02697$, and the threshold for appearance of $\chi^{+}\chi^{-}$ in the asymptotic state is $v_{*}\simeq 0.0505$. At small velocities, the Sommerfeld factor approaches to $\sim 1.715$ and velocity-independent, which is a typical behavior given by Yukawa potential. On the other hand, at large velocity, the perturbative picture becomes relevant and the Sommerfeld factor approaches to $1$. The behavior of the Sommerfeld factor near the threshold is rather complicated. Just above the threshold of $\chi^{+}\chi^{-}$, the on-shell charged Higgsinos state can be produced and the enhancement factor of $\chi^{+}\chi^{-}$ diverges as the inverse of relative velocity between them $1/v^{\chi^{\pm}}_{\rm{rel}}$. Such divergence is expected because at large distance, the Coulomb potential dominates the interactions. In this regime, an analytic estimate to the Sommerfeld factor of $\chi^{+}\chi^{-}$ is given by Landau and Lifshits (1991)
@@
Figure 1: The Sommerfeld enhancement factors for the ${}^{1}S_{0}$ partial-wave annihilation of $\chi_{1}^{0}\chi_{1}^{0}$, $\chi_{2}^{0}\chi_{2}^{0}$ and $\chi^{+}\chi^{-}$ channels. $v_{*}$ is the relative velocity between $\chi_{1}^{0}$s in the lightest two-particle state $\chi_{1}^{0}\chi_{1}^{0}$. For other heavier channels, the relative velocity $v_{\rm{rel}}$ is related to $v_{*}$ by the mass splittings.
@@ IV.1.2 Thermally averaged cross sections
$g_{\rm{eff}}=2+2(1+\Delta m_{\chi^{0}}/m_{\chi_{1}^{0}})^{3/2}e^{-x\Delta m_{\chi^{0}}/m_{\chi_{1}^{0}}}+4(1+\Delta m_{\chi^{\pm}}/m_{\chi_{1}^{0}})^{3/2}e^{-x\Delta m_{\chi^{\pm}}/m_{\chi_{1}^{0}}},$ (51)
@@
In Fig. 3, we plot $\langle\sigma v_{\rm{rel}}\rangle$ as a function of $x$ with and without including the Sommerfeld effect, adopting the mass-splitting parameters of model A. For the chosen values of mass splittings, the decoupling of heavier neutralino and charginos occurs at around $x\sim 10^{4}$. After the decoupling, the enhancement of $\langle\sigma v_{\rm{rel}}\rangle$ becomes constant, due to the velocity-independent Sommerfeld factor of $\chi_{1}^{0}\chi_{1}^{0}$ at very low velocities as shown in Fig. 1.
Figure 3: The perturbative and Sommerfeld enhanced thermally averaged cross sections $\langle\sigma v_{\rm{rel}}\rangle$ as functions of $x=m_{\chi_{1}^{0}}/T$, taking $m_{\chi^{0}_{1}}=0.8$, $1.1$, and $1.4\ {\rm TeV}$. Here, mass-splitting parameters of model A are adopted (i.e., $\Delta m_{\chi^{\pm}}=0.35\ {\rm GeV}$ and $\Delta m_{\chi^{0}}=0.1\ {\rm GeV}$).
@@ IV.2 Relic abundances
$\frac{dT}{dA}$ $=\left(1+\frac{T}{3h_{*}}\frac{dh_{*}}{dT}\right)^{-1}\biggl{[}-\frac{T}{A}+\frac{15T_{R}^{6}}{2\pi^{2}\sqrt{3}M_{\rm{Pl}}HT^{3}h_{*}A^{11/2}}\biggl{(}c_{\rho}^{1/2}A^{3/2}(1-\bar{B})\Phi$
@@ V Result
We also note here that, with our choices of $B_{\rm DM}$ parameter, the resultant relic density is insensitive to $B_{\rm DM}$ parameter; this is because the annihilation of the Higgsinos is efficient at the time of the reheating. Such a behavior is suggested by Eq. (38). As we have fixed $T_{R}$, changing $m_{\phi}$ only changes the number density of $\phi$ at the reheating. This effectively changes $B_{\rm DM}$ and the result is insensitive either.
@@ VI Conclusions and discussion
We have paid particular attention to the Higgsino LSP scenario, which is one of the popular and well-motivated scenario in the MSSM. We have performed a detailed analysis of the non-thermal production processes of Higgsino DM due to the decay of late-decaying scalar field. We have assumed that superparticles other than Higgsinos are so heavy that they do not affect the relic abundance of the Higgsino DM via coannihilation. We have numerically solved the Boltzmann equation governing the number density of Higgsino LSP, taking into account all the relevant annihilating channels among Higgsinos. Because Higgsinos have gauge quantum numbers and couple to electroweak gauge bosons, a non-perturbative effect, so called the Sommerfeld effect, may be significant; in our calculation, the Sommerfeld effect has been properly taken into account in the calculation of the annihilation cross sections of Higgsinos.
@@ A.0.2 $Q=0,S=1$
$\mathcal{H}_{Q=0}^{S=1}=\Phi_{0}^{V\dagger}\begin{pmatrix}2\Delta m_{\chi^{\pm}}-\Delta m_{\chi^{0}}-\frac{A^{2}\alpha_{2}e^{-m_{Z}r}}{4c_{w}^{2}r}-\frac{\alpha}{r}&-i\frac{\alpha_{2}e^{-m_{W}r}}{4r}\\ i\frac{\alpha_{2}e^{-m_{W}r}}{4r}&-\frac{\alpha_{2}e^{-m_{Z}r}}{4c_{w}^{2}r}\end{pmatrix}\Phi_{0}^{V}.$ (64)